OpenStudy (anonymous):

In a small bag of M&Ms, there are 10 orange candies and 4 blue candies. You select 1 candy, eat it, and then select another. Find P(you select first an orange candy and then a blue candy). Round your answer to the nearest hundredth.

OpenStudy (anonymous):

probability first one is orange is the ratio of orange pens to the total

OpenStudy (anonymous):

1/10?

OpenStudy (anonymous):

*candy

OpenStudy (anonymous):

slowly 10 orange candies and 4 blue candies how many candies are orange?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

right and how many are there total, orange and blue

OpenStudy (anonymous):

14

OpenStudy (anonymous):

14-1 (because you ate one)=13

OpenStudy (anonymous):

ok good so the probability you get an orange one on the first try is \[\frac{\text{number of orange}}{\text{total number of candies}}=\frac{10}{14}=\frac{5}{7}\]

OpenStudy (anonymous):

now that you have eaten an orange candy, how many candies are left in the bag?

OpenStudy (anonymous):

13 or 5/6

OpenStudy (anonymous):

oh i see, you already answered that as well, there are now 13 in the bag

OpenStudy (anonymous):

of those 13, how many are blue?

OpenStudy (anonymous):

Uhh 4

OpenStudy (anonymous):

right. so now the probability you pick a blue one is \[\frac{4}{13}\] take the two numbers, \[\frac{10}{14}\] and \[\frac{4}{13}\] and multiply them together to get your answer

OpenStudy (anonymous):

that is, compute \[\frac{5}{7}\times \frac{4}{13}\]

OpenStudy (anonymous):

Sorry my stupid computer died. So that would be 20/91

OpenStudy (anonymous):

wow you still here?

OpenStudy (anonymous):

yes it would be \(\frac{20}{91}\)

OpenStudy (anonymous):

this might have seemed like a lot of work, but really it was two short steps probability first one is orange is \(\frac{10}{14}\) probability that second is blue if you know the first one was orange is \(\frac{4}{13}\) and \[\frac{10}{14}\times \frac{4}{13}=\frac{20}{91}\]

OpenStudy (anonymous):

And if we rounded it to the nearest hundreth it would be 0.22!

OpenStudy (anonymous):

Alright what about this one.

OpenStudy (anonymous):

In a large bag of Skittles®, each of the 5 colors (red, orange, yellow, blue, green) occurs with the same probability. You reach in and select 2 candies.

OpenStudy (anonymous):

Find the probability that exactly one of the candies is orange.

OpenStudy (anonymous):

exactly one orange means one orange and one not orange right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

1/5?

OpenStudy (anonymous):

since there are 5 colors, and each has the same probability, the probability of each is \(\frac{1}{5}\) but that is not the answer to your question

OpenStudy (anonymous):

Hmm. So that means that they all have 1/5 chance of being selected but that isn't saying that one of them will be orange

OpenStudy (anonymous):

right lets say you pick two in succession, first pick and second pick you want the probability that exactly one of the picks is orange there are two ways to do this : { orange, not orange} or {not orange, orange}

OpenStudy (anonymous):

the probability of both of these is the same \[\frac{1}{5}\times \frac{4}{5}\] for each

OpenStudy (anonymous):

So i get 4/25?

OpenStudy (anonymous):

close \[\frac{4}{25}\] is the probability of each one of these events, but there are two of them, so you have to double it

OpenStudy (anonymous):

Double it.. so 16/50?

OpenStudy (anonymous):

i am not sure that any of this is making sense to you, but here is what is going on { orange AND not orange} for and you multiply and so it is \[\frac{1}{5}\times \frac{4}{5}\] then {not orange AND orange } again is \[\frac{4}{5}\times \frac{1}{4}\] since you want either one of these you ADD them and get \[\frac{4}{25}+\frac{4}{25}=\frac{8}{25}\]

OpenStudy (anonymous):

btw \[2\times \frac{4}{25}=\frac{8}{25}\] don't multiply top and bottom

OpenStudy (anonymous):

Ohhhh! See I was totally lost. Lol. So 8/25 ?

OpenStudy (anonymous):

yes i thought so it is \(\frac{8}{25}\) yes

OpenStudy (anonymous):

So if we divided it, it would be 0.32?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

yw